OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now? Numerical Answers Expected!

OpenStudy (ivettef365):

the way you set up the equation is as follows: Tim = x Johnson = x+10 and that is the same as Tim = x-3 Johnson = 2x

OpenStudy (ivettef365):

so now you do the equation as follows: x + x + 10 = x - 3 + 2x solve for x

OpenStudy (anonymous):

dont see it

OpenStudy (ivettef365):

do you know how to solve for x ??

OpenStudy (anonymous):

notreally

OpenStudy (ivettef365):

ok, you have x + x + 10 = 2x + 10 then you have x + 2x -3 = 3x -3

OpenStudy (anonymous):

ok

OpenStudy (ivettef365):

2x+10 = 3x-3, now combine like terms, when you switch you need to change signs as well so you have 10 + 3 = 3x - 2x

OpenStudy (ivettef365):

can you take it from there ??

OpenStudy (anonymous):

ya thanks

OpenStudy (ivettef365):

yw :)

OpenStudy (anonymous):

so its 13

OpenStudy (ivettef365):

good job

OpenStudy (anonymous):

Hey make sure you close the question if its been answered.. Thanks mate